Examine the function,
f(x)=2x^3- 9x^2- 60x+150
for extreme values.
1
Expert's answer
2020-03-04T18:46:09-0500

"y=2x^3-9x^2-60x+150\\
y'=6x^2-18x-60\\
6x^2-18x-60=0\\
x^2-3x-10=0"

According to Viet's theorem,

"x_1+x_2=3\\
x_1x_2=-10\\
x_1=5\\
x_2=-2"

"x\in(-\infty,-2), y'>0,\\
x\in(-2,5), y'<0,\\
x\in(5,\infty), y'>0."

Then

"x=-2" is a local maximum, "y_{max}=218" ;

"x=5" is a local minimum, "y_{min}=-125" .